OpenStudy (anonymous):

Solve for x. 4x^2+5x+24x+30=0

OpenStudy (anonymous):

grouping: 4x^2+24x+5x+30=0 4x(x+6)+5(x+6)=0 (4x+5)(x+6)=0 4x+5=0 or x+6=0 x= -5/4 or x=-6

OpenStudy (solomonzelman):

oksanaj is correct!

OpenStudy (anonymous):

Ah. I couldn't see the factors. Thank you!